# FlashBack : Efficient Retrieval-Augmented Language Modeling for Long Context Inference

### v3

@@ Results
In Figure 5, we compare the appending pattern (in-context), prepending pattern (in-context) and FlashBack in OPT models. We scale the model size from 125M to 6.7B. It is obvious that the acceleration of inference time is more effective in large models that have more layers and larger hidden sizes. In figure 6, we scale the max sequence length from 1088 to 3968 and observe significant improvement, as Llama 2 has a larger max sequence length, which is 4096 tokens. (Note: The test results for inference speed may vary depending on the hardware used and setting; our runtime tests are conducted on a single A100-80GB GPU setting to P0 state)
Figure 5: Time for OPT models generating a sequence of length 2048 tokens (including 128 tokens for retrieved content), given a input of 512 tokens, with retrieval stride of 16. We test appending and prepending patterns through in-context method for comparing with FlashBack. This is also a part of the ablation study.
@@ 4.2 Language Modeling
Figure 6: Time for Llama2-7B generating sequences with different length, given a input of 512 tokens, with retrieval stride of 16. This figure is for showing how the runtime increases in each distinct context pattern as sequence length scales up.
@@
We test FlashBack on Wikitext-2 (Merity et al., 2016) dataset, and three dataset: Arxiv, Freelaw and Stackexchange from the Pile (Gao et al., 2020). We report token-level average perplexity as the evaluation metric (excluding <MARK_L>,<MARK_R>). Our experiments use models of OPT (1.3B-6.7B) and GPT-2 (124M-1.5B). We compare FlashBack with the appending context pattern without fine-tuning and also the prepending context pattern. (For smaller models that are under 774M parameters, the results are in Appendix C Table 6)
@@ Results
Table 1 shows the results of different setup on the dev set. We can see that the appending context pattern applied directly on LLMs through the in-context method generates higher perplexity compared with the model using the prepending context pattern. Our result on fine-tuning with appending context patterns by using LoRA demonstrates that the model can adapt to such patterns after fine-tuning. Notably, with both the Marking Token and LoRA applied, the difference of perplexity results between two distinct context patterns becomes progressively alleviated as the size of the test model increases. And we also test GPT-2 from 124M to 1.5B to validate this observation. (See Table 6 in Appendix D)
4.3 Language Modeling with more retrieved content
@@
We use the same setup as section 4.2, except varying the number of retrieved content.
Perplexity
Model # Of Retrieved Content
1 2 4 8
GPT2-124M 25.00 24.06 24.65 25.72
GPT2-355M 16.05 15.58 15.91 16.50
GPT2-774M 13.28 12.94 13.21 13.67
GPT2-1.5B 12.49 12.19 12.43 12.83
OPT-125M 23.44 22.53 23.09 23.97
OPT-350M 16.89 16.31 16.67 17.29
OPT-1.3B 11.45 11.10 11.32 11.67
OPT-2.7B 10.48 10.14 10.32 10.63
OPT-6.7B 9.34 9.05 9.20 9.34
@@ Results
From the test results of varying the number of retrieved contents, the perplexity of all models we tested slightly worsens as the number of retrieved contents increases. We find the best perplexity comes with setting the number of retrieved contents to 2. We speculate the reasoning of this phenomena is that the second retrieved document is likely to be highly relevant with the first retrieved document. However, this experiment is only for demonstrating that FlashBack can be used for multi-documents retrieval. Further improvement on multi-documents retrieval is not our focus in this paper and it can be a direction in later work.
@@ 4.4 Implementation Details
We randomly sample from a discrete uniform distribution $U\{500,1000\}$ to get simulated token sequences to be the pseudo input and pseudo retrieved content. We maintain the length of simulated retrieved content to a fixed value of 128, then test models with different maximum sequence lengths. (Setup details are shown in Appendix Table 3). For The retrieval stride, all models are set to 16. All experiments are conducted by using a single GPU.
@@
For every $s$ tokens in the train set (refer to target), we sample from a discrete uniform distribution $U\{T/2,T-s\}$ to obtain the start position of the target, where $s$ is retrieval stride and $T$ is the maximum sequence length. For Wikitext-2, Wikitext-103, a superset of Wikitext-2, is used as a retrieval corpus. To avoid data leakage that target outputs are included in retrieved contents, we filter out retrieved content with titles that match Wikitext-2 during data construction. Owning to the much larger size of the Pile dataset, for Arxiv, Freelaw and Stackexchange, the validation split and 80% of the test split are mixtured as a retrieval corpus. And the rest of 20% of the test split are divided into two part equally as the training and the test dataset, respectively. We employ the BM25 retrieval and configure the query length to 16, indicating the utilization of the preceding 32 tokens from the target as input for the retriever. And we use the next token prediction objective for language modeling experiments. The loss is computed on target tokens in context.
@@
The trainable parameters for all models include two token embedding of <MARK_L> and <MARK_R>, and LoRA modules applied to all attention layers. The rank of weight matrices of LoRA is set to 16. The batch size is set to 16. The learning rates for GPT-2 models (124M-1.5B) are specified as [4e-4, 2e-4, 1e-4, 5e-5]. Correspondingly, the learning rates for OPT models (1.3B-6.7B) are specified as [1e-4, 5e-5, 5e-5]. We train for 5K additional steps, with 10% warm-up steps and a cosine learning rate schedule. For running on a single 24GB RAM GPU, it is feasible to use the bf16 format on the OPT-6.7B model and set the maximum sequence length to 512.
In our implementation, we uses the Transformers library (Wolf et al., 2020) and BM25 retrieval from the Pyserini library (Lin et al., 2021). And all fine-tuning experiments are conducted by using 4$\times$24GB RAM GPUs.
@@ 5 Conclusion
Prepending retrieved contents to the input without modification of model architecture has been a practicable method for LM to use external knowledge. However, the prepending context pattern hinders RALM from generalizing efficiently in considerable context length.
This paper presented the FlashBack method, enabling RALM with faster inference speed while keeping competitive model performance. We proved that the appending context pattern is more efficient than the prepending counterpart. Besides, our experiments demonstrated that LLMs can adapt to new context patterns using Marking Tokens with adaption fine-tuning (LoRA). And its performance in generation quality is closely equivalent to the LLM using prepending counterpart as size of the model increases and.
6 Limitations
In our experiments, the retrieval stride is a fixed value in each test, but it can be a dynamic variable, so there might be extra improvements. In run-time comparison experiments, due to constrained computational resources, we used simulated (pseudo) data for testing. It is because we did not find an appropriate public downstream task which fulfill following requirements: 1) long context, 2) multiple retrieval, 3) generated text becomes a new part of the context. However, we do believe this type of task exists in industrial applications. Additionally, we did not test our modeling on LLMs over size of 7B parameters since they generally require extensive computational resource.
Furthermore, in FlashBack, we did not scale the number of retrieved contents to a large value. It is practicable to generalize appending context patterns to more retrieved contents (for example, using the framework of REPLUG(Shi et al., 2023)).
@@ Appendix B Experiment Setup
Appendix C Ablation Study
Marking Token
We have already tested the model using an appending context pattern that is solely fine-tuned with LoRA without Marking Token to compare with FlashBack. In Table 1, test results of models with Marking Tokens and LoRA fine-tuning have lower PPL, which means fine-tuning the Marking Token could further boost the performance of the RALM and, therefore, meet our expectations. Further more, we also tested whether Marking Token and LoRA module add additional inference costs to our modeling. Our results on figure 5 show that FlashBack is slightly slower than model without these modules but the difference is tiny enough to be ignored compared to the overall increase of the inference speedup.
Retrieval Stride
Allocating the number of retrieval strides for each is a design choice that is related to model performance in both perplexity and inference time. As we mentioned above in section 3.1, In-Context RALM [Ram et al., 2023] choose to use a cherry-picked retrieval stride to balance the performance trade-off. They choose to use a relatively small stride value to have better generation quality (lower perplexity) without adding extensive inference cost. However, our experiments (see table 4 and 5 in appendix) on varying retrieval strides on FlashBack implied a different opinion. We find using 32 or 64 as retrieval stride does not degrade the model’s performance. When testing with a larger model such as OPT-6.7B, the test with a higher retrieval stride has the lowest perplexity compared to the test using same model with lower retrieval stride. Therefore, the trade-off between perplexity and inference time may not be a particularly evident phenomenon for some cases. As a consequence, we made a bold speculation that it might be an advantage for FlashBack for increasing the retrieval stride without intensively degrading the model’s generation quality in perplexity and still maintaining high-speed inference in long context scenarios. And this phenomenon may requires further demonstration in the future.
Perplexity
Model Retrieval Stride
16 32 64
GPT2-124M 28.11 23.64 21.80
GPT2-355M 16.46 15.66 15.80
GPT2-774M 13.46 13.12 13.24
GPT2-1.5B 12.59 12.28 12.24
OPT-125M 24.73 22.42 21.26
OPT-350M 17.29 16.71 16.71
OPT-1.3B 11.51 11.29 11.31
OPT-2.7B 10.70 10.38 10.23
OPT-6.7B 9.40 9.10 9.09

### v4

@@ Results
In Figure 4, we compare the appending pattern (in-context), prepending pattern (in-context), and FlashBack in OPT models. We scale the model size from 125M to 6.7B. Acceleration of inference time is more effective in large models that have more layers and larger hidden sizes. In Figure 5, we scale the maximum sequence length from 1088 to 3968 and observe a significant improvement, since Llama 2 has a larger maximum sequence length, which is 4096 tokens. (Note: The test results for inference speed may vary depending on the hardware used and setting; our runtime tests are conducted on a single A100-80GB GPU setting to P0 state)
Figure 4: Time for OPT models generating a sequence of length 2048 tokens (including 128 tokens for retrieved content), given a input of 512 tokens, with retrieval stride of 16.
@@ 4.2 Language Modeling
Figure 5: Time for Llama2-7B generating sequences with different length, given a input of 512 tokens, with retrieval stride of 16.
@@
We test FlashBack on Wikitext-2 (Merity et al., 2016) dataset and three datasets: Arxiv, Freelaw, and Stackexchange from the Pile (Gao et al., 2020). We report token-level average perplexity as the evaluation metric (excluding <MARK_L>,<MARK_R>).
Our experiments use models of OPT (1.3B-6.7B) and GPT-2 (124M-1.5B).
We compare FlashBack with the appending context pattern without fine-tuning and also the prepending context pattern. (For smaller models that are below 774M parameters, the results are in Appendix C Table 8)
@@ Results
Table 1 shows the results of different configurations in the dev set. We can see that the appending context pattern applied directly on LLMs through the in-context method generates higher perplexity compared to the model using the prepending context pattern. Our result on fine-tuning with appending context patterns by using LoRA demonstrates that the model can adapt to such patterns after fine-tuning. Notably, with both the Marking Token and LoRA applied, the difference of perplexity results between two distinct context patterns becomes progressively alleviated as the size of the test model increases. We also test GPT-2 from 124M to 1.5B to validate this observation. (See Table 8 in D). Furthermore, we test Llama 3.2-3B to validate our language modeling in one of the newest open source models (See Table 9 in E).
4.3 Language Modeling with Continuous Retrieval
@@
As the extra run-time cost of LoRA in comparison to the full-parameter is acceptable, we employ it in the following experiments. We continue to test FlashBack on the Wikitext-2 dataset (Merity et al., 2016) and design a multiple retrieval scenario. Specifically, given a token chunk $x_{1,...,n_{c}}$, we split it by retrieval stride to obtain prefixes: $x_{1,...,s},x_{1,...,2s},...,x_{1,...,n_{c}-s}$, where $n_{c}$ is the length of the chunk, and $s$ is the retrieval stride. Then we use each prefix with the corresponding retrieved passages to predict the next $s$ tokens. We report token-level average perplexity as the evaluation metric (excluding <MARK_L>,<MARK_R>).
@@ Results
Like the results on the language modeling task, a similar trend is observed in Table 2. Both Marking Token and LoRA are necessary to improve model performance. The performance gap between context patterns decreases as the model size increases.
[…]
OPT-1.3B P. 26.30
P. 14.82
P. 14.79
A. 15.41
A. 15.20
OPT-2.7B P. 22.73
P. 13.14
P. 13.14
A. 13.72
A. 13.49
OPT-6.7B P. 19.51
P. 11.65
P. 11.68
A. 12.18
A. 11.98
Table 2: Perplexity of OPT models on the language modeling with continuous retrieval task, evaluated on the validation set of WikiText-2. The retrieval stride is set to 16.
4.4 Question Answering
As improvement to the perplexity of the model results in performance improvement in downstream tasks, perplexity is a proxy metric for improved general-purpose abilities. In the In-Context RALM, retrieved passages boost model performance both on upstream and downstream tasks–language modeling and question answering, with perplexity and EM score as metrics, respectively. Although FlashBack can improve models on language modeling tasks, it is unclear whether it affects the downstream ability of models.
Setup
We test FlashBack in different settings (see Table 3) in Natural Questions (NQ) (Kwiatkowski et al., 2019) and TriviaQA(Joshi et al., 2017) open-domain question answering datasets to observe the effect of context pattern and the marking token on the QA task.
Results
Our results show that the context pattern does not evidently affect the model performance in the QA tasks after fine-tuning the model. However, the Marking Token can improve the model’s performance in QA tasks, especially for the test with the appending context pattern.
[…]
NQ TriviaQA
P. 17.0 43.3
P. 17.5 44.7
A. 18.7 42.5
A. 20.3 43.0
Table 3: QA Results of NQ, TriviaQA test set for OPT-6.7B. The model is fine-tuned on the language modeling with continuous retrieval task. We test with DPR retriever and keep top-2 retrieved documents.
@@ 4.5 Implementation Details
We randomly sample simulated token sequences from a discrete uniform distribution $U\{500,1000\}$ to become pseudo-input and pseudo-retrieved content. We maintain the length of simulated retrieved content to a fixed value of 128, then test models with different maximum sequence lengths. (Setup details are shown in the Appendix Table 5). For the retrieval stride, all models are set to 16. All run-time experiments are performed using a single NVIDIA A100 GPU.
@@
For every $s$ token in the train set (refer to target), we sample from a discrete uniform distribution $U\{T/2,T-s\}$ to obtain the start position of the target, where $s$ is the retrieval stride and $T$ is the maximum sequence length. For Wikitext-2, Wikitext-103, a superset of Wikitext-2, is used as a retrieval corpus. To avoid data leakage, target outputs are included in retrieved content, and we filter retrieved content with titles that match Wikitext-2 during data construction. Due to the much larger size of the Pile dataset, for Arxiv, Freelaw, and Stackexchange, the validation split and 80% of the test split are a mixture as a retrieval corpus. The rest of 20% of the test split is divided into two parts equally as the training and the test dataset, respectively. We employ BM25 retrieval and configure the query length to 16, indicating the use of the preceding 32 tokens of the target as input for the retriever. We use the next token prediction objective for language modeling experiments. The loss is computed on target tokens in context.
@@
The trainable parameters for all models include two token embeddings of <MARK_L> and <MARK_R>, and LoRA modules applied to all attention layers. The rank of LoRA weight matrices is set to 16. The batch size is set to 16. The learning rates for the GPT-2 models (124M-1.5B) are specified as [4e-4, 2e-4, 1e-4, 5e-5]. Correspondingly, the learning rates for the OPT models (1.3B-6.7B) are specified as [1e-4, 5e-5, 5e-5]. We train for 5K additional steps, with 10% warm-up steps and a cosine learning rate schedule. For running on a single 24GB RAM GPU, it is feasible to use the bf16 format on the OPT-6.7B model and set the maximum sequence length to 512.
In our implementation, we use the Transformers library (Wolf et al., 2020) and BM25 retrieval from the Pyserini library (Lin et al., 2021). All fine-tuning experiments are conducted by using 4$\times$24GB RAM GPUs.
5 Ablation Study
Marking Token
We have already tested the model using an appending context pattern that is solely fine-tuned with LoRA without Marking Token to compare with FlashBack. In Table 1, the test results of models with Marking Tokens and LoRA fine-tuning have lower PPL, which means that fine-tuning the Marking Token could further boost the performance of the RALM, and therefore meet our expectations. Furthermore, we also tested whether the Marking Token and LoRA modules add additional inference costs to our modeling. Our results in Figure 4 show that FlashBack is slightly slower than the model without these modules, but the difference is small enough to be ignored compared to the overall increase in the inference speed-up. In the QA test (Table 3), the results show that the marking token can be effective in predicting the model performance in both context patterns.
Perplexity
Model Retrieval Stride
16 32 64
GPT2-124M 28.11 23.64 21.80
GPT2-355M 16.46 15.66 15.80
GPT2-774M 13.46 13.12 13.24
GPT2-1.5B 12.59 12.28 12.24
OPT-125M 24.73 22.42 21.26
OPT-350M 17.29 16.71 16.71
OPT-1.3B 11.51 11.29 11.31
OPT-2.7B 10.70 10.38 10.23
OPT-6.7B 9.40 9.10 9.09
Retrieval Stride
Allocating the number of retrieval strides for each iteration is a design choice that is related to the performance of the model in both perplexity and inference time. As we mentioned above in section 3.1, the In-Context RALM (Ram et al., 2023) chooses to use a cherry-picked retrieval stride to balance the performance trade-off. They choose to use a relatively small stride value to have better generation (lower perplexity) without adding extensive inference cost. However, our experiments (see Table 4 and 6 in the appendix) on varying retrieval strides on FlashBack implied a different opinion. We find that using 32 or 64 as retrieval strides does not degrade the model’s performance. When testing with a larger model such as OPT-6.7B, the test with a higher retrieval stride has the lowest perplexity compared to the test using the same model with a lower retrieval stride. Therefore, the trade-off between perplexity and inference time may not be a particularly evident phenomenon for some cases. As a consequence, we made bold speculation that it might be an advantage for FlashBack to increase the retrieval stride without intensively degrading the model’s perplexity and still maintaining high-speed inference in long context input scenarios. This phenomenon may require further demonstration in the future.
@@ 6 Conclusion
Prepending retrieved contents to the input without modification of model architecture has been a practicable method for LM to use external knowledge. However, the prepending context pattern hinders RALM from efficiently generalizing at a considerable context length with a long input.
This paper presented the FlashBack method, which enables RALM with faster inference speed while maintaining competitive model performance. We proved that the appending context pattern is more efficient than the prepending counterpart. In addition, our experiments demonstrated that LLMs can adapt to new context patterns using Marking Tokens with adaption fine-tuning (LoRA). Its performance in generation quality is closely equivalent to the LLM using a prepending counterpart as the size of the model increases.
Limitations
FlashBack is specifically designed for RALM methods that use multiple retrievals by every $n$ token. Other works using retrieval strategies like retrieval once or retrieval by every $n$ text chunk are not covered in our paper and require further study. In our experiments, the retrieval stride is a fixed value in each test, but it can be a dynamic variable, so there might be additional improvements. In the run-time comparison experiment, due to constrained computational resources, we used simulated (pseudo) data for testing. This is because we did not find an appropriate public downstream task that fulfills the following requirements: 1) long context input, 2) multiple retrieval, and 3) generated text becomes a new part of the context. QA task generally comes with a short question, which is a short input, so the burden of RALM mainly lies on the retrieved documents, rather than the input. However, we believe that this type of scenario exists in some applications where long input is required with multiple retrievals. Additionally, we did not test our modeling on LLMs over the size of 7B parameters since they generally require extensive computational resources.
Furthermore, in FlashBack, we did not scale the number of retrieved contents to a large value. It is practicable to generalize appending context patterns to more retrieved contents (for example, using the REPLUG framework(Shi et al., 2023)).
Ethics Statement
This research explores the Retrieval-Augmented Generation (RAG) application leveraging Large Language Models (LLM). We call this method as Retrieval-Augmented Language Modeling (RALM).
Data in this research were obtained from publicly available datasets and sources that adhere to ethical guidelines, ensuring respect for user privacy and consent. Potential biases may exist inherently in both the external corpus and the LLMs. Using FlashBack for real applications requires additional precautions to avoid reinforcing stereotypes or propagating misinformation through the generation process.
@@ Appendix B Experiment Setup
Appendix C Language Modeling With More Retrieved Content
Setup
We use the same setup as in Section 4.2, except for varying the number of retrieved content.
Perplexity
Model # Of Retrieved Content
1 2 4 8
GPT2-124M 25.00 24.06 24.65 25.72
GPT2-355M 16.05 15.58 15.91 16.50
GPT2-774M 13.28 12.94 13.21 13.67
GPT2-1.5B 12.49 12.19 12.43 12.83
OPT-125M 23.44 22.53 23.09 23.97
OPT-350M 16.89 16.31 16.67 17.29
OPT-1.3B 11.45 11.10 11.32 11.67
OPT-2.7B 10.48 10.14 10.32 10.63
OPT-6.7B 9.34 9.05 9.20 9.34
Results
From the test results of varying the number of retrieved contents, the perplexity of all models we tested slightly worsens as the number of retrieved contents increases. We find that the best perplexity comes with setting the number of retrieved contents to 2. We speculate that the reasoning behind this phenomenon is that the second retrieved document is likely to be highly relevant to the first retrieved document. However, this experiment is only for demonstrating that FlashBack can be used for multi-document retrieval. Further improvement on multi-document retrieval is not our focus in this paper, and it can be a direction in later work.
@@ Appendix D Perplexity Results of GPT2 Model
We tested the GPT2 models on various model sizes in perplexity metrics. The results are presented in Table 8.
@@
Appendix E Perplexity Results of Llama 3.2
We tested the Llama 3.2-3B models in perplexity metrics to demonstrate FlashBack on one of the latest models. The experimental results are presented in table 9.
Model A./P. LoRA M.T. Wikitext-2 Arxiv Freelaw Stackexchange
Llama-3.2-3B N 9.02 5.33 6.16 6.14
P. 8.20 5.38 6.27 6.12
P. 6.92 4.93 5.95 5.74
P. 6.93 4.92 5.94 5.74
A. 33.00 32.03 38.63 38.06
A. 8.21 6.62 8.42 7.61
A. 7.56 5.40 6.71 6.25
Table 9: Perplexity of Llama-3.2-3B model on the entire validation set of WikiText-2 dataset and partial test set of Arxiv, Freelaw and Stackexchange datasets. The retrieval stride is set to 16.